Peanuts sell for 80.00 per kilogram and corn for 32.00 per kg. how many kg. of each should be mixed to form 150 kg of a mixture that can sell for 40.00 per kg.?
Answer by mananth(16946)

---------price ---------------- quantity
peanuts 80 ---------------- x kg
corn 32 --------------- 150 - x kg
Mixture 40.00% ---------------- 150

80 x + 32 ( 150 - x ) = 40 * 150

80x+4800-32x=6000
80x-32x =6000-4800
48x= 1200
/ 48
x = 25 kg $80.00 peanuts
125 kg $32.00 corn